For a certain stretched string tension=1225N consecutive resonance frequencies of 105,175,245 are observed. calculate wavelength in the funadamental mode. mass per unit length =0.25kg/m

Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello
v=root(T/mass per unit lenght)
105=35*3
175=35*5
245=35*7
so fundamental frequency =35
=>wavelength=v/f=root(1225/0.25)/35
Thanks & Regards
Arun Kumar
Btech IIT delhi
Askiitians Faculty